OpenStudy (acal21):

Find the degree 4 Taylor polynomial , T(sub4)(x) for the function f(x) = x^(3/2) centered at a =4

OpenStudy (acal21):

\[Find the degree 4 taylor polynomial, T_{4} for the function f(x) = x^{3/2} centered at a=4\]

OpenStudy (anonymous):

just gotta grind it till you find it \[a_0=f(4)=8\]\[a_1=f'(4)\] etc

OpenStudy (acal21):

ok so a_1=3 a_2=3/8 a_3=8 a_4=9/512

OpenStudy (acal21):

now what

OpenStudy (acal21):

well i solved it , but now i have to approximated the value of (3.9)^(3/2)

OpenStudy (acal21):

how would i do that in a calculator